Nearness of facilities Level of technology History of successful outsourcing The executives have determined four criteria weightings: Price, a weight of 0.1; Nearness, with 0.6; Technology, with 0.2 CRITERION Flexibility Trustworthi Price and History, with 0.1 a) U sing the factor-rating method, which country would you uble each of the weights used in part (a) (to 0.2, 1.2, 0.4, respectively). What effect does this have on your select? b) and 0.2, Delive answer? Why? Px 2.9 Ranga Ramasesh i firm that is trying to decide which one of four countries it s s the operations manager fora nager for a research for possible outsourcing providers. The first step is t criti- duction select a country based on cultural risk factors, which are cal to eventual business success with the provider. Ranga has reviewed outsourcing provider directories and found that thehis OM t four countries in the table that follows have an ample number nd found that the t of providers from which they can choose. To aid in the country selection step, he has enlisted the aid of a cultural expert, John Wang, who has provided ratings of the various criteria in the gs are on a 1 to 10 scale, where 1 is a low FACTO Labor risk and 10 is a high risk. Quality John has also determined six criteria weightings: Trust, with a weight of 0.4; Quality, with 0.2; Religious, with 0.1;Individualism,P with 0.1; Time, with 0.1; and Uncertainty, with 0.1. Using the Trust factor-rating method, which country should Ranga select? Px Logisti Price CULTURE SELECTION CRITERION Tech Man COSTA MEXICO PANAMA RICA PERU Trust Society value of quality work Religious attitudes Individualism attitudes Weig prov labo 7 on Time orientation attitudes is c 4 Uncertainty avoidance attitudes 4 2.10 Fernando Garzas firm wishes to use factor rating to
Answer #1

Mexico = 0.4*1+0.2*7+0.1*3+0.1*5+0.1*4+0.1*3 = 3.3
Panama = 0.4*2+0.2*10+0.1*3+0.1*2+0.1*6+0.1*2 = 4.1
Costa Rica = 0.4*2+0.2*9+0.1*3+0.1*4+0.1*7+0.1*4 = 4.4
Peru = 0.4*1+0.2*10+0.1*5+0.1*8+0.1*3+0.1*2 = 4.2

a. MEXICO has lowest total weighted score and is the best place for outsourcing.